OpenStudy (anonymous):

In a manufacturing plant that produces new computers, a 0.15 probability exists that a computer will be defective. If five computers are manufactured, what’s the probability that all of them will be defective?

OpenStudy (allank):

That should be the probability of one computer being defective*number of computers. 0.15*5

OpenStudy (anonymous):

so you multiply by the percent ?

OpenStudy (allank):

Yep.

OpenStudy (anonymous):

i got 0.75

OpenStudy (allank):

Lemme think about the logic of that..

OpenStudy (anonymous):

should i round it to become .76

OpenStudy (allank):

Sorry about this change. Actually, it won't be 0.15*5, but: (probability of a computer being defective)^(no of computers)=(0.15)*(0.15)*(0.15)...=0.15^5

OpenStudy (allank):

So the probability is actually 7.59E-5 percent.